If X and Y are the roots of the equation ax2+bx+c=0 , then the roots of the equation ax2-bx(x-1)+c(x-1)2=0 are :
 

Grade:11

1 Answers

Charchit Tailong
8 years ago
X+Y=-b/a
XY=c/a
from 2nd equation
ax2-bx2+b+cx2+c-2cx
(a-b+c)x2-2cx+b=0
alpha and beta are 2 roots
alpha+beta=2c/(a-b+c)
alpha*beta=b/(a-b+c)
put X+Y and X*Y in this